# Revisiting and Maximizing Temporal Knowledge in Semi-supervised Semantic Segmentation

### v1

@@ 3.2 Previous Guidance: Revisiting Temporal Knowledge
In contrast to existing methods, we directly utilize previous models and generate pseudo-label guidance from their predictions without relying on a complex pipeline of dual EMA-based methods or a filtering mechanism. In particular, we store multiple models at different epochs that meet the specified criteria during training and use their pseudo-labels as additional guidance, referred to as previous guidance. As training progresses, the decoupling between the student and previously saved teachers increases, allowing for the acquisition of different prediction views. However, as decoupling becomes more pronounced, the positive effects of self-training from correct pseudo-labels may diminish due to the use of outdated teachers. Therefore, previous guidance is used in conjunction with the standard guidance obtained from the current student. In addition, we define the maximum length ($N$) of the list for storing previous models and replace the oldest teacher with a newer one when this limit is exceeded to avoid using excessively outdated teachers. Formally, one previous model is randomly selected from the previous model list $\{T_{1},T_{2},...,T_{N}\}$. Then, this model processes the same weakly perturbed image, $x^{w}$, used in the standard flow and produces predictions $p^{w^{\prime}}$. The previous guidance is obtained by: $y^{w^{\prime}}\text{= argmax}(p^{w^{\prime}})$, and it supervises three predictions according to Eq. (5). In this way, we produce diverse pseudo-labels, as in prior studies [30, 31, 28, 29], by leveraging different previous models encompassing varied perspectives of temporal knowledge, without complex additional components or heavy computational burden.
@@
Save Criteria. In this approach, storing the appropriate previous models is crucial for generating diverse and reliable pseudo-labels. Some studies [62, 63, 21] that employ intermediate models for ensembling or filtering in image recognition and segmentation save the model at regular intervals (e.g., every 20 epochs in a total of 100 epochs). By contrast, we save the model when it achieves the best performance on the validation set to ensure the stability of the previous guidance. The weights and performances of the neural network can vary significantly during the optimization process, and excessively large fluctuations may increase negative impacts. These negative impacts can be exacerbated by label scarcity and a pronounced class imbalance in semi-supervised semantic segmentation compared to standard image recognition. In addition, the periodic saving method requires additional hyperparameter searches to determine appropriate intervals. Therefore, we adopt our saving approach to achieve stability and simplicity (refer to Table IX for related experiments).
The efficacy of the previous guidance can be intuitively explained as follows. When using only standard guidance, two training scenarios arise based on whether the prediction is correct or incorrect. In cases where the prediction is incorrect, the network is trained in the wrong direction. In contrast, given the standard and previous guidance, four scenarios can be considered based on whether they are correct or incorrect (standard-previous): (1) correct-correct, (2) correct-incorrect, (3) incorrect-correct, and (4) incorrect-incorrect. Through case (3), the network receives an additional opportunity to be guided in the right direction, away from the wrong one. Although model training may be hindered through case (2), leading to significant fluctuations, we empirically demonstrate that the positive effects of the proposed method outweigh the negative effects (which is further supported by the analysis of training stability in Section 4.4.1). In addition, previous guidance can help mitigate the network’s catastrophic forgetting problem [64, 65], where previously learned knowledge is forgotten when acquiring new knowledge. In particular, this phenomenon can be more pronounced in self-training and class-imbalance scenarios (i.e., in our scenario) due to the lack of labeled data, potentially leading to significant performance fluctuations in poorly behaved classes [19]. In semi-supervised [20] and representation learning [61], some studies have demonstrated that utilizing temporal knowledge helps mitigate the catastrophic forgetting problem and stabilizes training. Based on this fact, we explore the learning stability of the proposed method in Section 4.4.1.
@@ 3.3 Maximizing Efficacy of Previous Guidance
One way to improve the reliability of the predictions involves using network ensemble techniques. These techniques have been widely used in various domains as a promising method for improving performance. In particular, several studies [62, 63] in the field of image recognition have demonstrated that ensembles of multiple intermediate models obtained during training also improve prediction accuracy and diversity. Therefore, we utilize network ensembling to improve the reliability of the previous guidance. In designing this method, we also consider computational complexity and pseudo-label diversity. To this end, given the list that includes $N$ previous models, we randomly select $K$ ($K\leq N$) models for each iteration to mitigate the increase in computational complexity while ensuring pseudo-label diversity. However, this approach, which always ensembles $K$ models, may not guarantee the diversity of pseudo-labels, as noted in Dual Teacher [31]. This problem can worsen for large values of $K$.
Therefore, to provide reliable and diverse pseudo-labels to the student network, we propose a highly randomized ensemble strategy comprising the following two ideas:
  • Random Selection: For each iteration, we randomly select a varying number of teachers, $k$, ranging from 1 to $K$. For example, $k$=1,2, or 3 can be selected for each iteration when $K$=3. With this approach, selecting a large $k$ tends to yield consistent pseudo-labels, enhancing reliability, while a small $k$ contributes to increased diversity in the pseudo-labels. In addition, computational costs are lower than those incurred when using a fixed number $K$. This strategy enables the student network to obtain stable and diverse guidance, which functions as a robust regularizer, aiding network optimization.
@@
In this way, the proposed randomized ensemble strategy can improve the reliability and diversity of the previous guidance while mitigating the increase in computational complexity.
@@ 4 Experiments
4.1 Experimental Configuration
4.1.1 Datasets
PASCAL VOC [66] is a widely used benchmark dataset in semantic segmentation, comprising 21 distinct categories (20 object types and a background class). This dataset contains 10,582 images designated for training purposes, segmented into two subsets based on annotation quality: 1,464 images constitute the high-quality subset, characterized by detailed annotations, whereas the remaining 9,118 images form the coarse subset, featuring less detailed annotations. This study follows established protocols for a fair comparison [28, 67, 32, 39]. In particular, three training protocols are considered based on the criteria for selecting labeled images.
  • Original: A protocol whereby labeled images are exclusively sourced from the high-quality subset.
  • Blended: A protocol that entails a random selection of labeled images from the total dataset.
  • Priority: A protocol where the selection of labeled images is first derived from the high-quality subset; if not sufficient, it is complemented by additional images from the coarse subset.
Cityscapes dataset [68], tailored for the semantic analysis of urban street scenes, comprises 2,975 high-resolution images for training and 500 images for validation, primarily focusing on 19 categories within urban environments. Consistent with previous studies [21, 32, 29], we evaluate this dataset using various label partitions, specifically 1/16, 1/8, 1/4, and 1/2 of the total number of labels.
COCO dataset [69], notable for its complexity and scale, comprises 118k training and 5k validation images. This dataset features dense annotations across 81 classes set in various indoor and outdoor scenarios. Considering the performance plateaus observed in datasets such as Pascal and Cityscapes and the higher number of classes, the COCO dataset has emerged as a practical and valuable benchmark for evaluating advanced algorithms in semi-supervised segmentation. Following existing studies [32], we validate the proposed method using 1/256, 1/128, and 1/64 label partitions.
ADE20K dataset [70], containing more diverse scenes and a greater number of classes (150 categories) than COCO, comprises 20,210 images for training and 2,000 images for validation. Using the label partitions from existing studies [33], we evaluate the proposed method on 1/128, 1/64, and 1/32 partitions.
In all protocols, training images not selected as labeled images are utilized as unlabeled images.
4.1.2 Architecture and Implementation Details
Consistent with existing literature, we employ ResNet-50 and ResNet-101 backbones [71] as encoders for the Pascal VOC, Cityscapes, and ADE20K datasets. For the COCO dataset, Xception-65 [72] is adopted as the backbone. We use DeepLabV3+ [3] as the segmentation head and set its output stride to 16 for efficient training.
@@ 4.1.2 Architecture and Implementation Details
For the training setups, each mini-batch comprises 8 labeled and 8 unlabeled images. The proposed method is trained for 80, 240, 30, and 40 epochs for the PASCAL, Cityscapes, COCO, and ADE20K datasets, respectively, using the SGD optimizer. The learning rates are initially set to 0.001, 0.005, 0.004, and 0.004 for these datasets, respectively, and are managed using a polynomial learning rate scheduler. Moreover, we use 321/513, 801, 513, and 513 random crops for these datasets, respectively. For image augmentation, we use common weak (e.g., resize, crop, and flip) and strong (e.g., color transformations, grayscale, cutmix, and blur) data augmentations, as in UniMatch [32]. For the hyperparameters of standard consistency flow, $\tau$ is set to 0 for Cityscapes and 0.95 for the other datasets. The dropout rate for feature perturbation is set to 0.5.
@@
For the hyperparameters of PrevMatch, $\tau$ is set to 0.9 for Pascal, 0 for Cityscapes, and 0.95 for COCO and ADE20k. The maximum length ($N$) of the previous list is set to eight for Pascal and Cityscapes and five for COCO and ADE20K. The upper bound number $K$ for the random selection is set to three. For weight $\lambda$, as previous guidance should correct the model before it becomes overfitted in the wrong direction (i.e., confirmation bias [17]) during the middle of training, setting an appropriate weight $\lambda$ is crucial. Additionally, as the initial model typically exhibits poor performance, we implement a warmup schedule for the weight of the proposed flow, similar to the commonly used polynomial learning rate scheduler that includes a warmup phase. Experiments are conducted in the following environments: UBUNTU 20.04, Python 3.10.4, PyTorch 1.12.1, CUDA 11.3, and NVIDIA RTX 3090Ti or A6000 GPUs. The training is performed using one GPU for Pascal and two GPUs for the other datasets.
TABLE I: Comparison with state-of-the-art methods on the Original protocol of Pascal VOC dataset. The numeric values in the header (e.g., 92) represent the number of labeled images used for training. All methods are trained using ResNet-50/101 and DeepLabV3+. The number of trainable whole networks and input resolution are reported. The values of PrevMatch are averaged over three runs. The evaluation metric is the mean IoU (%).
Pascal [Original set] Encoder #Trainable Networks Resolution # Labeled images (Total: 10582)
92 183 366 732 1464
Supervised Baseline R-50 $\times$1 5132 44.0 52.3 61.7 66.7 72.9
PseudoSeg [35] [ICLR’21] R-50 $\times$1 5132 54.9 61.9 64.9 70.4 71.0
PC2Seg [73] [ICCV’21] R-50 $\times$1 5132 56.9 64.6 67.6 70.9 72.3
CPCL [60] [TIP’23] R-50 $\times$2 5122 61.9 67.0 72.1 74.3 -
AugSeg [39] [CVPR’23] R-50 $\times$1 5122 64.2 72.2 76.2 77.4 78.8
UniMatch [32] [CVPR’23] R-50 $\times$1 3212 71.9 72.5 76.0 77.4 78.7
Dual Teacher [31] [NeurIPS’23] R-50 $\times$1 3212 70.8 74.5 76.4 77.7 78.2
PrevMatch (ours) R-50 $\times$1 3212 73.4 75.4 77.5 78.6 79.3
Supervised Baseline R-101 $\times$1 5132 45.1 55.3 64.8 69.7 73.5
CPS [28] [CVPR’21] R-101 $\times$2 5122 64.1 67.4 71.7 75.9 -
ReCo [74] [ICLR’22] R-101 $\times$1 3212 64.8 72.0 73.1 74.7 -
PS-MT [30] [CVPR’22] R-101 $\times$1 5122 65.8 69.6 76.6 78.4 80.0
ST++ [21] [CVPR’22] R-101 $\times$1 3212 65.2 71.0 74.6 77.3 79.1
U2PL [67] [CVPR’22] R-101 $\times$1 5122 68.0 69.2 73.7 76.2 79.5
GTA-Seg [75] [NeurIPS’22] R-101 $\times$2 5132 70.0 73.2 75.6 78.4 80.5
PCR [57] [NeurIPS’22] R-101 $\times$1 5132 70.1 74.7 77.2 78.5 80.7
DGCL [76] [CVPR’23] R-101 $\times$1 5132 70.5 77.1 78.7 79.2 81.6
CCVC [47] [CVPR’23] R-101 $\times$2 5122 70.2 74.4 77.4 79.1 80.5
iMAS [40] [CVPR’23] R-101 $\times$1 5132 68.8 74.4 78.5 79.5 81.2
AugSeg [39] [CVPR’23] R-101 $\times$1 5122 71.1 75.5 78.8 80.3 81.4
UniMatch [32] [CVPR’23] R-101 $\times$1 3212 75.2 77.2 78.8 79.9 81.2
ESL [58] [ICCV’23] R-101 $\times$1 5132 71.0 74.1 78.1 79.5 81.8
LogicDiag [59] [ICCV’23] R-101 $\times$1 5132 73.3 76.7 77.9 79.4 -
LogicDiag + MKD [59] [ICCV’23] R-101 $\times$2 5132 74.7 77.2 78.4 80.1 -
Diverse Co-T. (2-cps) [29] [ICCV’23] R-101 $\times$2 3212 74.8 77.6 79.5 80.3 81.7
Diverse Co-T. (3-cps) [29] [ICCV’23] R-101 $\times$3 3212 75.4 76.8 79.6 80.4 81.6
PrevMatch (ours) R-101 $\times$1 3212 77.0 78.5 79.6 80.4 81.6
TABLE II: Comparison with state-of-the-art methods on the Blended protocol of Pascal VOC dataset. All methods are trained using ResNet-50 and DeepLabV3+. The fractional values indicate the ratio of labeled images used for training.
Pascal [Blended set] Resolution 1/16 1/8 1/4
Supervised Baseline 5132 62.4 68.2 72.3
Mean Teacher [26] [NeurIPS’17] 5122 66.8 70.8 73.2
CCT [44] [CVPR’20] 5122 65.2 70.9 73.4
GCT [55] [ECCV’20] 5122 64.1 70.5 73.5
CutMix-Seg [38] [BMVC’20] 5122 68.9 70.7 72.5
CAC [77] [CVPR’21] 3202 70.1 72.4 74.0
CPS [28] [CVPR’21] 5122 72.0 73.7 74.9
UCC [46] [CVPR’22] 5122 74.1 74.8 76.4
PS-MT [30] [CVPR’22] 5122 72.8 75.7 76.4
UniMatch [32] [CVPR’23] 3212 74.5 75.8 76.1
AugSeg [39] [CVPR’23] 5122 74.7 76.0 77.2
iMAS [40] [CVPR’23] 5132 74.8 76.5 77.0
CCVC [47] [CVPR’23] 5122 74.5 76.1 76.4
PrevMatch (ours) 3212 75.6 76.4 76.3
5132 76.0 77.1 77.6
@@
TABLE III: Comparison with state-of-the-art methods on the Priority protocol of Pascal VOC dataset. All methods are trained using ResNet-101 and DeepLabV3+.
Pascal [Priority set] Resolution 1/16 1/8 1/4
Supervised Baseline 5132 70.6 75.0 76.5
U2PL [67] [CVPR’22] 5122 77.2 79.0 79.3
UniMatch [32] [CVPR’23] 5132 80.9 81.9 80.4
AugSeg [39] [CVPR’23] 5122 79.3 81.5 80.5
Dual Teacher [31] [NeurIPS’23] 5132 80.1 81.5 80.5
PrevMatch (ours) 5132 81.4 81.9 80.8

### v2

@@ 3.2 Previous Guidance
To address the coupling problem between student and teacher networks, we directly utilize previous models and generate pseudo-label guidance from their predictions, instead of using Mean Teacher or co-training approaches. In particular, we store multiple models at different epochs that meet the specified criteria during training and use their pseudo-labels as additional guidance, referred to as previous guidance. As training progresses, the decoupling between the student and previously saved teachers increases, allowing for the acquisition of different prediction views. However, as decoupling becomes more pronounced, the positive effects of self-training from correct pseudo-labels may diminish due to the use of outdated teachers. Therefore, previous guidance is used in conjunction with the standard guidance obtained from the current student. In addition, we define the maximum length ($N$) of the list for storing previous models and replace the oldest teacher with a newer one when this limit is exceeded to avoid using excessively outdated teachers.
Formally, one previous model is randomly selected from the previous model list $\{T_{1},T_{2},...,T_{N}\}$. Then, this model processes the same weakly perturbed image, $x^{w}$, used in the standard flow and produces predictions $p^{w^{\prime}}$. The previous guidance is obtained by: $y^{w^{\prime}}\text{= argmax}(p^{w^{\prime}})$, and it supervises the strongly perturbed prediction according to Eq. (4). In this way, we produce diverse pseudo-labels, as in prior studies [liu2022perturbed, na2023switching, chen2021semi, li2023diverse], by leveraging different previous models encompassing varied perspectives of temporal knowledge, without complex additional components or heavy computational burden.
3.2.1 Save Criteria.
@@
In this approach, storing the appropriate previous models is crucial for generating diverse and reliable pseudo-labels. Some studies [huang2017snapshot, wang2022efficient] that employ intermediate models for ensembling in image recognition save the model at regular intervals (e.g., every 20 epochs in a total of 100 epochs). In contrast to standard image recognition, the weights and performances of the neural network can vary significantly during the optimization process in semi-supervised segmentation (i.e., our case) due to label scarcity and a pronounced class imbalance. Therefore, we save the model when it achieves the best performance on the validation set to ensure the stability of the previous guidance. In addition, this approach does not require additional hyperparameter searches, such as determining appropriate intervals. Therefore, we adopt our saving approach to achieve stability and simplicity.
The efficacy of the previous guidance can be intuitively explained as follows. When using only standard guidance, two training scenarios arise based on whether the prediction is correct or incorrect. In cases where the prediction is incorrect, the network is trained in the wrong direction. In contrast, given the standard and previous guidance, four scenarios can be considered based on whether they are correct or incorrect (standard-previous): (1) correct-correct, (2) correct-incorrect, (3) incorrect-correct, and (4) incorrect-incorrect. Through case (3), the network receives an additional opportunity to be guided in the right direction, away from the wrong one (Refer to Sec. 8.2 in the supplementary material for further discussion).
@@ 3.3 Maximizing Efficacy of Previous Guidance
One way to improve the reliability of the predictions involves using network ensemble techniques. These techniques have been widely used in various domains as a promising method for improving performance. In particular, several studies [huang2017snapshot, wang2022efficient] in the field of image recognition have demonstrated that ensembles of multiple intermediate models obtained during training also improve prediction accuracy and diversity. Therefore, we utilize network ensembling to improve the reliability of the previous guidance. In designing this method, we also consider computational complexity and pseudo-label diversity. To this end, given the list that includes $N$ previous models, we randomly select $K$ ($K\leq N$) models for each iteration to mitigate the increase in computational complexity while ensuring pseudo-label diversity. However, this approach, which always ensembles $K$ models, may not guarantee the diversity of pseudo-labels, as noted in Dual Teacher [na2023switching]. This problem can worsen for large values of $K$. Therefore, to provide reliable and diverse pseudo-labels to the student network, we propose a highly randomized ensemble strategy comprising the following two ideas:
  • Random Selection: For each iteration, we randomly select a varying number of teachers, $k$, ranging from 1 to $K$. For example, kk=1,2, or 3 can be selected for each iteration when KK=3. With this approach, selecting a large $k$ tends to yield consistent pseudo-labels, enhancing reliability, while a small $k$ contributes to increased diversity in the pseudo-labels. In addition, computational costs are lower than those incurred when using a fixed number $K$. This strategy enables the student network to obtain stable and diverse guidance, which functions as a robust regularizer, aiding network optimization.
@@
In this way, the proposed randomized ensemble strategy probabilistically generates pseudo-labels with varying degrees of reliability and diversity across the training process, thereby enabling broader and more flexible exploration of the pseudo-label space.
@@ 4 Experiments
4.1 Datasets
(1) Pascal VOC dataset [everingham2010pascal] contains 10,582 training images, segmented into two subsets based on annotation quality: 1,464 images constitute the high-quality subset, whereas the remaining 9,118 images form the coarse subset. In particular, two training protocols are considered based on the criteria for selecting labeled images: Original protocol whereby labeled images are exclusively sourced from the high-quality subset. Blended protocol that entails a random selection of labeled images from the total dataset. (2) Cityscapes dataset [cordts2016cityscapes], tailored for the semantic analysis of urban street scenes, comprises 2,975 high-resolution images for training and 500 images for validation, primarily focusing on 19 categories within urban environments. (3) ADE20K dataset [zhou2017scene], containing more diverse scenes and a greater number of classes (150 categories) than the COCO dataset [lin2014microsoft], comprises 20,210 images for training and 2,000 images for validation. In all protocols, training images not selected as labeled images are utilized as unlabeled images.
4.2 Implementation Details
@@
We select various methods that potentially exhibit confirmation bias and coupling problems as baselines, which use the input perturbation method based on a single network. For the four baseline methods (FixMatch, UniMatch, AugSeg, and MPMC), we employ ResNet-50/101 [he2016deep] as the encoder and DeepLabV3+ [chen2018encoder] as the segmentation head. For UniMatchV2, we use the DINOv2 [oquab2023dinov2] encoder and the DPT [ranftl2021vision] segmentation model. Specifically, UniMatch and UniMatchV2 are the primary baselines in our study. For these two baselines, each mini-batch comprises 8 labeled and 8 unlabeled images. Furthermore, they are trained for 80, 240/180, and 40/60 epochs for the PASCAL, Cityscapes, and ADE20K datasets, respectively. The learning rates are initially set to 1e-3/5e-6, 5e-3/5e-6, and 4e-3/5e-6 for these datasets, respectively, and are managed using a polynomial learning rate scheduler. Moreover, we use 321/518, 801/798, and 513/518 random crops for these datasets, respectively. For image augmentation, we use common weak (e.g., resize, crop, and flip) and strong (e.g., color transformations, grayscale, cutmix, and blur) data augmentations, as in UniMatch [yang2023revisiting]. For the hyperparameters of standard consistency flow, $\tau$ is set to 0 for Cityscapes and 0.95 for the other datasets.
@@
For the hyperparameters of PrevMatch, $\tau$ is set to 0.9 for Pascal, 0 for Cityscapes, and 0.95 for ADE20k. The maximum length ($N$) of the previous list is set to eight for Pascal and Cityscapes and six for ADE20K. The upper bound number $K$ for the random selection is set to three. For weight $\lambda$, as previous guidance should correct the model before it becomes overfitted in the wrong direction (i.e., confirmation bias [arazo2020pseudo]) during the middle of training, setting an appropriate weight $\lambda$ is crucial. Additionally, as the initial model typically exhibits poor performance, we implement a warmup schedule for the weight of the proposed flow, similar to the commonly used learning rate decay scheduling that includes a warmup phase. Refer to Sec. 7.6 in the supplementary material for further details.
Pascal [Original set] Encoder 92 183 366 732 1464
AugSeg [zhao2023augmentation] [CVPR’23] RN-50 64.2 72.2 76.2 77.4 78.8
Dual Teacher [na2023switching] [NeurIPS’23] RN-50 70.8 74.5 76.4 77.7 78.2
FixMatch [sohn2020fixmatch] [NeurIPS’20] RN-50 63.8 70.3 73.2 76.8 77.8
+ PrevMatch RN-50 66.1 (+2.3) 73.1 (+2.8) 75.7 (+2.5) 77.8 (+1.0) 78.5 (+0.7)
UniMatch [yang2023revisiting] [CVPR’23] RN-50 71.9 72.5 76.0 77.4 78.7
+ PrevMatch RN-50 73.4 (+1.5) 75.4 (+2.9) 77.5 (+1.5) 78.6 (+1.2) 79.3 (+0.6)
U2PL [wang2022semi] [CVPR’22] RN-101 68.0 69.2 73.7 76.2 79.5
PS-MT [liu2022perturbed] [CVPR’22] RN-101 65.8 69.6 76.6 78.4 80.0
Diverse Co-T.(3cps) [li2023diverse] [ICCV’23] RN-101 75.4 76.8 79.6 80.4 81.6
CorrMatch [sun2024corrmatch] [CVPR’24] RN-101 76.4 78.5 79.4 80.6 81.8
U2PL+ [wang2024using] [IJCV’24] RN-101 69.3 73.4 75.0 77.1 79.5
AugSeg [zhao2023augmentation] [CVPR’23] RN-101 71.1 75.5 78.8 80.3 81.4
+ PrevMatch RN-101 73.4 (+2.3) 77.3 (+1.8) 80.0 (+1.2) 80.9 (+0.6) 82.0 (+0.6)
UniMatch [yang2023revisiting] [CVPR’23] RN-101 75.2 77.2 78.8 79.9 81.2
+ PrevMatch RN-101 77.0 (+1.8) 78.5 (+1.3) 79.6 (+0.8) 80.4 (+0.5) 81.6 (+0.4)
AllSpark [wang2024allspark] [CVPR’24] MiT-B5 76.1 78.4 79.8 80.8 82.1
SemiVL [hoyer2023semivl] [ECCV’24] CLIP-B 84.0 85.6 86.0 86.7 87.3
UniMatchV2-S [yang2025unimatch] [TPAMI’25] DINOv2-S 79.0 85.5 85.9 86.7 87.8
+ PrevMatch DINOv2-S 84.4 (+5.4) 86.2 (+0.7) 86.7 (+0.8) 87.5 (+0.8) 88.2 (+0.4)
UniMatchV2-B [yang2025unimatch] [TPAMI’25] DINOv2-B 86.3 87.9 88.9 90.0 90.8
+ PrevMatch DINOv2-B 87.2 (+0.9) 88.6 (+0.7) 90.5 (+1.6) 90.6 (+0.6) 91.3 (+0.5)
@@
Table 1: Comparison with state-of-the-art methods on the Original protocol of Pascal VOC. The evaluation metric is the mean IoU (%).
Pascal [Blended set] Encoder 1/16 1/8 1/4
Mean Teacher [tarvainen2017mean] RN-50 66.8 70.8 73.2
PS-MT [liu2022perturbed] RN-50 72.8 75.7 76.4
CorrMatch [sun2024corrmatch] RN-101 77.6 77.8 78.3
AllSpark [wang2024allspark] MiT-B5 78.3 80.0 80.4
FixMatch [sohn2020fixmatch] RN-50 72.5 73.0 74.2
+ PrevMatch RN-50 73.9 74.3 75.1
Gain ($\Delta$) (+1.4) (+1.3) (+0.9)
UniMatch [yang2023revisiting] RN-101 76.5 77.0 77.2
+ PrevMatch RN-101 77.8 78.0 78.0
Gain ($\Delta$) (+1.3) (+1.0) (+0.8)
UniMatchV2-S [yang2025unimatch] DINOv2-S 82.2 83.6 83.8
+ PrevMatch DINOv2-S 83.3 84.5 84.5
Gain ($\Delta$) (+1.1) (+0.9) (+0.7)
Table 2: Comparison with state-of-the-art methods on the Blended protocol of Pascal VOC dataset.